OpenStudy (anonymous):

Doris, Eloise, and Fran are sisters. Doris is the oldest, and Fran is the youngest. The sum of the ages of the three sisters is 56. Doris is 4 more than twice Fran's age. The difference between Eloise and Fran's ages is 9 years. Write a system of equations to determine the ages. A. Step 1: define your variables B. Step 2: write THREE equations

OpenStudy (anonymous):

@jim_thompson5910 please help.

jimthompson5910 (jim_thompson5910):

how far did you get

OpenStudy (anonymous):

D+E+F=56 That is all I know how to do.

jimthompson5910 (jim_thompson5910):

so for "A. Step 1: define your variables" you should have D = current age of Dorris E = current age of Eloise F = current age of Fran

OpenStudy (anonymous):

I did that.

OpenStudy (anonymous):

I just don't know how to write the other 2 equations.

jimthompson5910 (jim_thompson5910):

`The sum of the ages of the three sisters is 56` you got `D+E+F=56` which is correct

jimthompson5910 (jim_thompson5910):

`Doris is 4 more than twice Fran's age` which equation can we form?

OpenStudy (anonymous):

D= 4+2x ?

jimthompson5910 (jim_thompson5910):

but remember that F = frans age

OpenStudy (anonymous):

D=4+2F

jimthompson5910 (jim_thompson5910):

yeah or D = 2F + 4

jimthompson5910 (jim_thompson5910):

`The difference between Eloise and Fran's ages is 9 years` what equation forms?

OpenStudy (anonymous):

E-F=9

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Thank you! Can you please help through several other problems? I just looked in my binder and I have more problems.

jimthompson5910 (jim_thompson5910):

I'll help with 3 more